Given triangle ABC whose sides AB=15 in, AC=25, BC=30. From point D on side AB, a line DE is drawn to point AC such that angle ADE is equal to angle ABC. If perimeter of ADE is 28 in, find the length of segments BD and CE
1
Expert's answer
2020-03-19T16:08:30-0400

since DE and BC are parallel, triangles ABC and ADE are similar.


The ratio of sides of ADE to ABC = 28/70 = 2/5


So,

CE = AC * 3/5 = 15

BD = AB * 3/5 = 9